# The Negative Participation Paradox in Three-Candidate Instant Runoff Elections

### v1

@@ 3.2. Impartial Culture
    the limiting probability that an election demonstrates a negative participation paradox is approximately $0.151084\approx 15.11\%.$.
@@
    the limiting probability that an election demonstrates a negative participation paradox is $0.617525\approx 61.75\%$, assuming the IRV winner is not the plurality winner.
@@
The proof of Proposition 3 is similar to that of Proposition 2. It can be found in the Appendix. We prove Proposition 2 below.
Proof.
Proof of (i)
@@
Step 1 First, we find the probability that $A$ and $B$ are the plurality and IRV winners respectively. Following the arguments of [26], the probability that (1) and (3) are satisfied is equal to the area of the spherical simplex $S$ defined by these inequalities on the surface of the unit sphere in $\mathbb{R}^{3}$, divided by the area of this sphere. Let
@@
$\alpha_{13}$ $=\angle(\mathbf{v}_{1},\mathbf{v}_{3})=\cos^{-1}(\frac{-\mathbf{v}_{1}\cdot\mathbf{v}_{3}}{\lVert\mathbf{v}_{1}\rVert\lVert\mathbf{v}_{3}\rVert})=\cos^{-1}(\frac{4}{2\cdot\sqrt{6}})=\cos^{-1}(\frac{1}{\sqrt{6}})$
@@
Step 2: Next, we find the probability that additionally (4) is satisfied. Following [10], we introduce the modified inequality with parameter $t\in[0,1],$
(7) $x_{1}+x_{2}>x_{3}+x_{4}+t(x_{4}-x_{3}).$
@@
$dVol_{n}=\frac{1}{n-1}\sum_{1\leq j<k\leq n}Vol_{n-2}(S_{j}\cap S_{k})d\alpha_{jk}.$
@@
$Vol(S_{1}\cap C_{4})$ $=\angle(P_{124},P_{134})=\cos^{-1}(\frac{0}{\sqrt{2}\cdot 2\sqrt{3}})=\frac{\pi}{2}$
$Vol(S_{2}\cap C_{4})$ $=\angle(P_{124},P_{234})=\cos^{-1}(\frac{\sqrt{3}t}{\sqrt{5t^{2}+3}})$
$Vol(S_{3}\cap C_{4})$ $=\angle(P_{134},P_{234})=\cos^{-1}(\frac{\sqrt{2}t}{\sqrt{5t^{2}+3}})$
@@
$\frac{1}{2\pi^{2}}\left[0.04077671(2\pi^{2})+\frac{1}{2}(-0.99679+0.11231+0.268772)\right]=0.02518064.$
@@
Since the choice of $A$ and $B$ as plurality and IRV winners was arbitrary, the probability of a negative participation paradox is $6\times 0.02518064\approx 0.151084$
@@
Proof of (ii) Since a negative participation paradox can only occur when the IRV and plurality winners are distinct, the probability that an election demonstrates a negative participation paradox given that these winners are different is equal to the probability of negative participation paradox divided by the probability that the IRV and plurality winners are distinct. The latter was determined in Step 1; hence this probability is equal to $0.02518386/0.04077671\approx 0.617525$
3.3. Analysis of IAC and IC Results
Overall, a negative participation paradox is twice as likely under the IC than the IAC model, and somewhat more likely assuming the IRV and plurality winners are distinct. This is to be expected, as the IAC model frequently produces preference profiles containing a majority candidate while this is impossible under IC in the limiting case. Under both models, the probability of a negative participation paradox is slightly higher for complete voter preferences than for partial voter preferences; the relationship is reverse when conditioned on the elections having distinct IRV and plurality winners. This is possibly because the probability that these winners are distinct is somewhat lower for partial preferences than for complete preferences (in the IC case, 21% versus 24%), With partial ballots, it is less likely that a candidate ranking second in first-place votes will gain sufficient additional votes, after the plurality loser is eliminated, to overcome the vote total of the plurality winner. We compare these results with an empirical analysis in the next section.
4. Empirical Results
In this section we present empirical results using a large database of real-world IRV elections. Our data comes from three sources. First, we use single-winner political elections from the United States with at least three (not write-in) candidates. This data is available at the FairVote data repository [23], and contains ballot data from municipal elections in cities such as San Francisco, CA and Minneapolis, MN. The repository also contains ballot data from federal elections in Alaska and Maine. Second, we use single-winner elections from the American Psychological Association, which generally is willing to share ballot data for ranked-choice presidential elections and elections for the Board of Directors. Some of this data is available on preflib.org [14], and some was shared directly with the first author. Third, we use single-winner political elections from Scotland from a database of Scottish local government elections. The ballot data for these elections is available at https://github.com/mggg/scot-elex.
In total, we have access to the preference profiles for 361 single-winner IRV elections with at least three candidates. Many of these elections contain more than three candidates; because we focus on the three-candidate case, for each election we run the IRV algorithm until three candidates remain and use the resulting profile. As with Example 1, any paradox we find in the resulting three-candidate profile will also be demonstrated in the original election. The reason is that when we create the ballots which produce the paradox only the first and last rankings matter; the intermediate rankings can be filled in arbitrarily using the remaining candidates from the original candidate set. Of course, it is possible that an election with four or more candidates demonstrates a paradox which cannot be found when we reduce down to three candidates; our methodology would not find such elections.
In Section 3 we analyze the likelihood of negative participation paradoxes in two cases: when all voters provide complete preferences and when some voters cast partial ballots. To provide comparisons to those results, with our real data we also give results using fully complete preferences and using partial ballots. In real elections voters often provide partial rankings of the candidates, and thus our results from partial ballots represent the “actual” results, the results obtained from using the actual ballot data. To create complete preferences, we proportionally fill in partial ballots using the available complete ballots to create hypothetical results for the case that all voters submit complete preferences.
To see how we fill in the ballots, consider the preferences profile from Table 1. We need to take the bullet votes for $J$ and extend them to complete ballots over the three candidates. Since $1430+1012=2442$ voters rank $J$ first and provide complete rankings, we extend $1430/2442*100\%$ of the 1149 bullet votes for $J$ to ballots of the form $J\succ S\succ W$ and $1012/2442*100\%$ of the bullet votes for $J$ to ballots of the form $J\succ W\succ S$, rounding to the nearest integer. As a result, 673 of the 1149 bullet votes are extended to $J\succ S\succ W$ and 476 are extended to $J\succ W\succ S$. This methodology is used on every three candidate profile to create complete preferences.
Jurisdiction USA APA Scotland
Elections 304 27 30
No Majority Cand 191 23 30
IRV Winner $\neq$ plurality winner 18, 30 3, 3 5, 8
Paradoxes 7, 17 3, 3 2, 3
Table 4. Results from real-world elections. In the final two rows, the first number is obtained using the actual ballots and the second is obtained from ballots which have been completed.
Our results are summarized in Table 4. The top row shows the number of elections from each jurisdiction and the second row shows the number of elections without a majority candidate. The third row shows the number of elections in which the IRV winner is different from the plurality winner. The first number is the number of such elections when using the actual data with partial ballots; the second number is the number of such elections when using complete preferences, created as described above. The fourth row shows the number of participation paradoxes, using the same format. For example, there are 191 American political elections without a majority candidate. Of these, the IRV winner is not the plurality winner in 18 (respectively 30) elections when using the actual ballot data (respectively, complete preference data). Seven American elections demonstrate a paradox when using the actual data, which rises to 17 when we complete the ballots.
Election Num. Voters Ballot Range
2020 San Francisco Board of Supervisors D7 38321 798-1605
2021 Minneapolis City Council Ward 2 8907 80-224
2023 Minneapolis City Council Ward 8 7899 2973-4090
2010 Oakland Mayor 113217 2315-3951
2016 Oakland School Director D5 12950 3151-4267
2022 Oakland School Director D4 26432 38-598
2021 New York City Rep Primary D5 8182 773-830
2005 APA President 14079 2275-2780
2007 APA President 12925 2397-2717
2020 APA Board of Directors Race 2 6227 1093-1883
2021 Argyll Bute By-Election Isle of Bute Ward 1804 26-38
2021 Highland By-Election Aird Ward 3321 266-266
Table 5. Each election which demonstrates a negative participation paradox when using the actual ballot data.
Table 5 provides a complete list of elections which demonstrate a paradox when using the actual ballots. The table also provides the range of the number of ballots which can be added to produce the paradox. For example, the third row shows that we can add anywhere from 2973 to 4090 ballots of the form $W\succ J\succ S$ to create the paradox in Example 1. Note that the last election in the table just barely produces a paradox, as the paradox can be demonstrated using only 266 ballots exactly.
When using the actual ballot data, in total we find 12 elections which demonstrate this paradox. From Table 4, we obtain an empirical probability of $12/361=3.3\%$ that the paradox occurs. When we condition on the non-existence of a majority candidate we obtain a probability of $12/244=4.9\%$, which rises to $12/26=46.2\%$ when we restrict to elections when the IRV and plurality winners are different. The corresponding three probabilities for complete preferences are 6.4%, 9.4%, and 56.1%, respectively. As with the IAC and IC models we obtain higher probabilities for complete preferences for the two conditional probabilities. Unlike those models, we obtain higher probabilities for complete preferences for the unconditioned probabilities as well. Overall, the frequency with which the paradox of negative participation occurs in the database is lower than that which is predicted by the IAC and IC models, as expected since these models tend to provide upper bounds for the probabilities of paradoxical behavior. It is interesting, though, that some of the empirical probabilities are not significantly lower than the theoretical probabilities. Often, an empirical estimate of the probability that an election demonstrates a particular voting paradox is much lower than what is predicted by a model like IAC or IC. When we assume the IRV and plurality winners are different, the empirical probability for negative participation paradoxes is surprisingly close to the probabilities reported in Section 3.
5. Comparison to other paradoxes
[…]
% Voters $p_{A}$ $p_{B}$ $p_{C}$
1st choice $A$ $B$ $C$
2nd choice $B$ $A$ $B$
3rd choice $C$ $C$ $A$
Table 6. A table illustrating that an election can demonstrate a negative participation paradox even if the plurality loser controls a small percentage of the first-place votes.
Furthermore, to create such a paradox we can control the type and number of ballots added to the original ballot data. Thus, we expect that such paradoxes are exhibited more frequently in real-world data than other paradoxes. Prior empirical research bears this out. Using essentially the same database of American elections that we use, [6] found four elections which demonstrate an upward monotonicity paradox, three which demonstrate a downward monotonicity paradox, and one which demonstrates a no-show paradox. The authors did not include the APA elections in that study, but they evaluated those elections using the same methodology as in [6] and found no paradoxes of any kind. The study in [16] included the 30 Scottish elections we evaluate as part of a much larger study of single transferable vote elections, and of these 30 elections found only one which demonstrated a paradox (this election demonstrated both an upward monotonicity paradox and a no-show paradox). In total, of the 361 elections in our database there are five documented elections which demonstrate an upward monotonicity paradox, three which demonstrate a downward monotonicity paradox, and two which demonstrate a no-show paradox. As in this paper, all of the previously documented paradoxes occurred at the “three candidate level” in the sense that the paradox can be demonstrated after eliminating all but the final three candidates.
6. Conclusion
We conclude with two final comments. First, IRV proponents sometimes claim that a strength of IRV is that the method can choose a different (and presumably more deserving) winner than the plurality winner. This is a natural point to make, since if the two winners always coincide then there is no reason to use IRV, and arguably the plurality winner sometimes does not deserve to win an election (e.g. if the plurality winner is a Condorcet loser). However, in some sense our work shows that this feature of IRV comes with a price: when the IRV and plurality winners differ, it is often the case that the election demonstrates a negative participation paradox. That is, when IRV “works properly” by choosing a more deserving candidate than the plurality winner, often the plurality winner could have been the IRV winner if only the turnout of some voters who don’t like the plurality winner were increased.
Second, we make no claim regarding the normative status of negative participation paradoxes. Social choice theorists care about paradoxes such as monotonicity and participation paradoxes in part because such paradoxes show there is some mathematical irrationality built into the mechanics of IRV. Put simply, sometimes IRV does not behave rationally in response to changes in the ballot data, and this seems normatively undesirable. On the other hand, paradoxes in this vein are hypothetical–negative participation paradoxes especially so. In Example 2, it is extremely unlikely that there were thousands of potential voters in Minneapolis’ eighth ward whose favorite candidate was White and whose least favorite was Stevenson, and these voters abstained from the actual election. Example 2 is built using thousands of people who probably do not exist, which perhaps ultimately says nothing about whether IRV is a “good” or “appropriate” voting method.
Regardless of one’s normative stance on these paradoxes, the results in this paper contribute to the literature on voting paradoxes by providing limiting probabilities regarding negative participation paradoxes in the three-candidate case under the impartial culture model, empirical probabilities using a large database of real-world elections, and limiting probabilities in the three-candidate case for some cases left over from [11] under the impartial anonymous culture model.
Appendix
@@ 6. Conclusion
Step 2: Next, we find the probability that additionally (6) is satisfied. We introduce the inequality $x_{1}+x_{2}+x_{3}>x_{4}+x_{5}+x_{6}+t(x_{6}-x_{5})$. When $t=0$ this reduces to a (2); when $t=1$ we get the required extra condition.
@@
$Vol(S_{1}\cap C_{4})$ $=\angle(P_{124},P_{134})=\cos^{-1}(\frac{0}{\sqrt{2}\cdot 2\sqrt{18}})=\frac{\pi}{2}$
$Vol(S_{2}\cap C_{4})$ $=\angle(P_{124},P_{234})=\cos^{-1}(\frac{18t}{\sqrt{2}\cdot 3\sqrt{2(13t^{2}+9)}})=\cos^{-1}(\frac{3t}{\sqrt{13t^{2}+9}})$
$Vol(S_{3}\cap C_{4})$ $=\angle(P_{134},P_{234})=\cos^{-1}(\frac{18t}{3\sqrt{2}\cdot 3\sqrt{2(13t^{2}+9)}})=\cos^{-1}(\frac{t}{\sqrt{13t^{2}+9}})$
@@
Proof of (ii) As with the case with complete ballots, the probability that an election demonstrates a negative participation paradox given that these winners are different is equal to the probability of negative participation paradox divided by the probability that the IRV and plurality winners are distinct. The latter was determined in Step 1; hence this probability is equal to $0.023976912/0.03563737\approx 0.672803$

### v3

@@ 3.2 Impartial Culture
    the limiting probability that an election demonstrates a negative participation paradox is approximately $0.157737\approx 15.77\%$.
@@
    the limiting probability that an election demonstrates a negative participation paradox is $0.644717\approx 64.47\%$, assuming the IRV winner is not the plurality winner.
@@
The proofs of both propositions can be found in the Appendix.
3.3 Analysis of IAC and IC Results
Overall, a negative participation paradox is twice as likely under the IC than the IAC model, and much more likely assuming the IRV and plurality winners are distinct. This is to be expected, as the IAC model frequently produces preference profiles containing a majority candidate while this is impossible under IC in the limiting case. Under the IC model, the probability when using complete preferences is slightly higher than when using partial preferences; we investigate why this should also be the case in real-world elections in Section 4. Under IAC, the difference between partial and complete preferences for the unconditioned probability is negligible.
Under the IAC model, the probability that an election does not contain a majority candidate is $7/16=43.75\%$ in the complete preference case and $145/256=56.64\%$ in the partial ballot case. Thus, even though the unconditioned limiting probabilities are approximately equal, the probability conditioned on the absence of a majority candidate is lower in the partial ballot case.
Under both models, the probability for the case of partial preferences is larger than for complete preferences when we assume the IRV and plurality winners are distinct. This is possibly because the probability that these winners are different is somewhat lower for partial preferences than for complete preferences (in the IC case, 21% versus 24%). With partial ballots it is less likely that a candidate ranking second in first-place votes will gain sufficient additional votes, after the plurality loser is eliminated, to overcome the vote total of the plurality winner.
Refer to caption Refer to caption
Figure 1: The ballot ranges of elections which demonstrates a paradox under the IAC and IC models with 3901 voters.
In addition to calculating probabilities under each model, we also examine the range of the possible number of ballots (given by Inequalities 4 and 8) which can be added to demonstrate a paradox. If an election exhibits a paradox, we say the ballot range of the election is the maximum number of ballots which can be added to demonstrate the paradox minus the minimum number of ballots. In Example 2, we can add as many as 3389 ballots to demonstrate the paradox and as few as 2924, and thus the ballot range for that election is $3389-2924=465$. The range of possible number of ballots gives a measure of how “close” the election is to one that does not demonstrate the paradox—the larger the range, the more susceptible the election.
To investigate such ballot ranges, for each model we generated 100,000 elections with complete ballots using 3901 voters. For each election which demonstrated a negative participation paradox we calculated the ballot range; the results are displayed in Figure 1 for elections with complete ballots. We also generated box plots for elections with partial ballots but the corresponding plots are not markedly different. The left box plot for IAC uses the 7268 generated elections which demonstrated a paradox under this model; the box plot for IC uses 15128 elections. [^3] Note that 15128 out of 100,000 runs gives an estimated probability a bit far from the probability of 15.77% given in Proposition 3. The reason is that even with a large electorate of 3901 voters, approximately 2400 simulated elections contain at least two candidates with the same number of first-place votes. If we ignore such elections with ties (which is the case for a limiting probability) then our estimated probability becomes much closer to 15.77%. The median ballot range under the IC model is 32 and the maximum is 182. Thus, relative to an electorate size of 3901, under IC elections which generate a paradox tend to have small ballot ranges. In particular, 227 of the generated elections which demonstrate a paradox have a ballot range of zero, and the paradoxical outcome is as “close” as possible to an election which does not exhibit a paradox. Under the IAC model the median ballot range is 806 and the maximum is 1934. Thus, while IAC produces far fewer paradoxes than IC, this model tends to produce much larger ballot ranges. In this case, only two of the generated elections produce a ballot range of zero.
We compare our IC and IAC results with an empirical analysis in the next section.
4 Empirical Results
In this section we present empirical results using a large database of real-world IRV elections. Our data comes from three sources. First, we use single-winner political elections from the United States with at least three (not write-in) candidates. This data is available at the FairVote data repository (Otis, 2022), and contains ballot data from municipal elections in cities such as San Francisco, CA and Minneapolis, MN. The repository also contains ballot data from federal elections in Alaska and Maine. Second, we use single-winner elections from the American Psychological Association, which generally is willing to share ballot data for ranked-choice presidential elections and elections for the Board of Directors. Some of this data is available on preflib.org (Mattei & Walsh, 2013), and some was shared directly with the first author. Third, we use single-winner political elections from Scotland from a database of Scottish local government elections. The ballot data for these elections is available at https://github.com/mggg/scot-elex.
In total, we have access to the preference profiles for 361 single-winner IRV elections with at least three candidates. Many of these elections contain more than three candidates; because we focus on the three-candidate case, for each election we run the IRV algorithm until three candidates remain and use the resulting profile. As with Example 1, any paradox we find in the resulting three-candidate profile will also be demonstrated in the original election. The reason is that when we create the ballots which produce the paradox only the first and last rankings matter; the intermediate rankings can be filled in arbitrarily using the remaining candidates from the original candidate set. Of course, it is possible that an election with four or more candidates demonstrates a paradox which cannot be found when we reduce down to three candidates; our methodology would not find such elections. We note also that in trying to characterize the likelihood of such a paradox in real settings, we are assuming that voters’ relative preferences over the remaining three candidates in each election would not change had those candidates been the only ones running. Such an assumption seems reasonable given the rationale of the IRV algorithm itself, which sequentially eliminates candidates and bumps up remaining candidates in each voter’s preference order.
In Section 3 we analyzed the likelihood of negative participation paradoxes in two cases: when all voters provide complete preferences and when some voters cast partial ballots. To provide a similar empirical analysis, with our real data we also give results using fully complete preferences and using partial ballots. In real elections voters often provide partial rankings of the candidates, and thus our results from partial ballots represent the “actual” results, the results obtained from using the actual ballot data. To create complete preferences, we proportionally fill in partial ballots using the available complete ballots to create hypothetical results for the case that all voters submit complete preferences.
To see how we fill in the ballots, consider the preferences profile from Table 1. We need to take the bullet votes for $J$ and extend them to complete ballots over the three candidates. Since $1430+1012=2442$ voters rank $J$ first and provide complete rankings, we extend $1430/2442*100\%$ of the 1149 bullet votes for $J$ to ballots of the form $J\succ S\succ W$ and $1012/2442*100\%$ of the bullet votes for $J$ to ballots of the form $J\succ W\succ S$, rounding to the nearest integer. As a result, 673 of the 1149 bullet votes are extended to $J\succ S\succ W$ and 476 are extended to $J\succ W\succ S$. This methodology is used on every three-candidate profile to create complete preferences.
In Section 3 we do not create complete preference data proportionally in this manner; instead, we generate complete ballots separately from partial ballots. The reason for this difference in methodology is that when using theoretical models like IAC or IC, we have no pre-existing data from which to draw, and if we treat the cases separately then we can use the geometric techniques outlined in Section 3 to calculate probabilities for both cases. If we were to create complete ballots proportionally after first using IAC or IC to generate partial ballots, the complete ballot probabilities are similar but not equal to the probabilities in Propositions 1 and 3. (We confirmed this using simulations.) Thus, a comparison between our theoretical and empirical results for complete ballots is less useful than a comparison between our results for partial ballots, although both comparisons are still valuable.
Jurisdiction USA APA Scotland Total
Elections 304 27 30 361
No Majority Cand 191 23 30 244
IRV Winner $\neq$ Plurality Winner 18, 30 3, 3 5, 8 26, 41
Paradoxes 7, 17 3, 3 2, 3 12, 23
Table 4: Results from real-world elections. In the final two rows, the first number is obtained using the actual ballots and the second is obtained from ballots which have been completed.
Partial (actual) Completed
Probability of a paradox 3.3% 6.4%
Given No Majority Cand 4.9% 9.4%
Given IRV Winner $\neq$ Plurality Winner 46.2% 56.1%
Table 5: Probability of a negative participation paradox in real-world elections.
Our results are summarized in Table 4. The top row shows the number of elections from each jurisdiction and the second row shows the number of elections without a majority candidate. The third row shows the number of elections in which the IRV winner is different from the plurality winner. The first number is the number of such elections when using the actual data with partial ballots; the second number is the number of such elections when using complete preferences, created as described above. The fourth row shows the number of negative participation paradoxes, using the same format. For example, there are 191 American political elections without a majority candidate. Of these, the IRV winner is not the plurality winner in 18 (respectively 30) elections when using the actual ballot data (respectively, complete preference data). Seven American elections demonstrate a paradox when using the actual data, which rises to 17 when we complete the ballots. By summing all the instances of the paradox across all the election categories and dividing by the appropriate denominator we obtain the probabilities obtained in Table 5.
As with the IAC and IC models, we obtain higher probabilities for complete preferences for the two conditional probabilities. Unlike those models, we obtain higher probabilities for complete preferences for the unconditioned probabilities as well. This is partly due to the fact that completing ballots can “introduce” paradoxes to elections where no potential existed in the actual voter profile.
We illustrate this phenomenon using the 2022 IRV mayoral election from Corvallis, Oregon.
Example 3.
The 2022 election for mayor of Corvallis, Oregon, marked the first use of IRV for a mayoral election in the city. Ignoring write-in candidates, the election contained the three candidates Hogg ($H$), Maughan ($M$), and Struthers ($S$). The preference profile for the actual election is displayed in the top of Table 6.
In the actual election, the initial vote totals are 5535, 7917, 7147 for $H$, $M$, and $S$, respectively. After the elimination of Hogg, Maughan defeats Struthers 9893 to 9859. Note that the plurality and IRV winners agree, and thus the election cannot demonstrate a negative participation paradox.
If we proportionally complete the partial ballots then we obtain the profile in the bottom of Table 6. In this modified election Hogg is still eliminated first, but Struthers defeats Maughan 10349 votes to 10250 in the last round. Note that Inequalities (3) and (5) are satisfied, and if we add $z$ ballots of the form $H\succ S\succ M$ where $z\in[2383,4534]$ then Maughan becomes the winner.
Num. Voters 847 1976 2712 372 2380 5165 517 2398 4232
1st choice $H$ $H$ $H$ $M$ $M$ $M$ $S$ $S$ $S$
2nd choice $M$ $S$ $H$ $S$ $H$ $M$
3rd choice $S$ $M$ $S$ $H$ $M$ $H$
Num. Voters 2333 3202 2497 5420 2585 4562
1st choice $H$ $H$ $M$ $M$ $S$ $S$
2nd choice $M$ $S$ $H$ $S$ $H$ $M$
3rd choice $S$ $M$ $S$ $H$ $M$ $H$
Table 6: The actual preference profile for the 2022 Corvallis mayoral election (Top) and the profile with preferences completed proportionally (Bottom).
Overall, the frequency with which the paradox of negative participation occurs in the real-world database is lower than that which is predicted by the IAC and IC models, which is to be expected since these models tend to provide upper bounds for the probabilities of paradoxical behavior. It is interesting, though, that some of the empirical probabilities are not significantly lower than the theoretical probabilities. Often, an empirical estimate of the probability that an election demonstrates a particular voting paradox is much lower than what is predicted by a model like IAC or IC. When we assume the IRV and plurality winners are different, the empirical probability for negative participation paradoxes is surprisingly close to the probabilities reported in Section 3.
We complete our analysis by examining the numbers of ballots that can be added to create a negative participation paradox in the twelve elections that were susceptible to the paradox using the partial (actual) ballots. Table 7 provides a complete list of these elections along with the total number of voters, the number of ballots that can be added to produce the paradox (expressed as a ballot interval), and the ballot range divided by the original number of voters. For example, the third row shows that we can add anywhere from 2924 to 3389 ballots of the form $W\succ J\succ S$ to create the paradox in Example 1, corresponding to a range of $3389-2924=465$. Dividing 465 by 7899 yields 5.9%. Note that the last election in the table just barely produces a paradox, as the paradox can be demonstrated using only 266 ballots exactly. Recall that under the IC model, the median ballot range for generated elections demonstrating the paradox was 32 with an electorate size of 3901, which corresponds to a percentage of 0.08%. Under the IAC model, the median range was 806, yielding 20.7%. The ranges found in the actual elections, therefore, lie somewhere between these two extremes.
Election Num. Voters Ballot Interval Range Perc
2020 San Francisco Board of Supervisors District 7 38321 798-1606 2.1%
2021 Minneapolis City Council Ward 2 8907 80-224 1.6%
2023 Minneapolis City Council Ward 8 7899 2924-3389 5.9%
2010 Oakland Mayor 113217 2315-3951 1.4%
2016 Oakland School Director District 5 12950 2876-4325 11.2%
2022 Oakland School Director District 4 26432 38-598 2.1%
2021 New York City Rep Primary District 50 8182 773-830 0.7%
2005 APA President 14079 2275-2780 3.6%
2007 APA President 12925 2397-2718 2.5%
2020 APA Board of Directors Race 2 6227 1093-1883 12.7%
2021 Argyll Bute By-Election Isle of Bute Ward 1804 26-38 0.7%
2021 Highland By-Election Aird Ward 3321 266-266 0.0%
Table 7: Each election which demonstrates a negative participation paradox when using the actual ballot data. The $\dagger$ denotes that these two elections were previously found (using different methodology) in (McCune, 2024).
5 Comparison to other paradoxes
[…]
% Voters $p_{A}$ $p_{B}$ $p_{C}$
1st choice $A$ $B$ $C$
2nd choice $B$ $A$ $B$
3rd choice $C$ $C$ $A$
Table 8: A table illustrating that an election can demonstrate a negative participation paradox even if the plurality loser controls a small percentage of the first-place votes.
Furthermore, to create such a paradox we can control the type and number of ballots added to the original ballot data. Thus, we expect that such paradoxes are exhibited more frequently in real-world data than other paradoxes. Prior empirical research bears this out. Using essentially the same database of American elections that we use, (Graham-Squire & McCune, 2023) found four elections which demonstrate an upward monotonicity paradox, three which demonstrate a downward monotonicity paradox, and one which demonstrates a no-show paradox. The authors did not include the APA elections in that study, but they evaluated the APA elections using the same methodology as in (Graham-Squire & McCune, 2023) and found no paradoxes of any kind. The study in (McCune & Graham-Squire, 2023) included the 30 Scottish elections we evaluate as part of a much larger study of single transferable vote elections, and of these 30 elections found only one which demonstrated a paradox (this election demonstrated both an upward monotonicity paradox and a no-show paradox). In total, of the 361 elections in our database there are five documented elections which demonstrate an upward monotonicity paradox, three which demonstrate a downward monotonicity paradox, and two which demonstrate a no-show paradox. As in this paper, all of the previously documented paradoxes occurred at the “three candidate level” in the sense that the paradox can be demonstrated after eliminating all but the final three candidates.
6 Conclusion
We conclude with two final comments. First, IRV proponents sometimes claim that a strength of IRV is that the method can choose a different (and presumably more deserving) winner than the plurality winner. This is a natural point to make, since if the two winners always coincide then there is no reason to use IRV, and arguably the plurality winner sometimes should not win an election (e.g. if the plurality winner is a Condorcet loser). However, in some sense our work shows that this feature of IRV comes with a price: when the IRV and plurality winners differ, it is often the case that the election demonstrates a negative participation paradox. That is, when IRV “works properly” by choosing a more deserving candidate than the plurality winner, often the plurality winner could have been the IRV winner if only the turnout of some voters who don’t support the plurality winner were increased.
Second, we make no claim regarding the normative status of negative participation paradoxes. Social choice theorists care about paradoxes such as monotonicity and no-show paradoxes in part because such paradoxes show there is some mathematical irrationality built into the mechanics of IRV. Put simply, sometimes IRV does not behave rationally in response to changes in the ballot data, and this seems normatively undesirable. On the other hand, paradoxes in this vein are hypothetical–negative participation paradoxes especially so. In Example 2, it is extremely unlikely that there were thousands of potential voters in Minneapolis’ eighth ward whose favorite candidate was White and whose least favorite was Stevenson, and these voters abstained from the actual election. Example 2 is built using thousands of people who probably do not exist, which perhaps ultimately says nothing about whether IRV is a “good” or “appropriate” voting method. Furthermore, while negative participation paradoxes may be normatively undesirable, IRV may have other positive features which justify its use, despite the susceptibility of IRV to various voting paradoxes.
Regardless of one’s normative stance on these paradoxes, the results in this paper contribute to the literature on voting paradoxes by providing limiting probabilities regarding negative participation paradoxes in the three-candidate case under the impartial culture model, empirical probabilities using a large database of real-world elections, and limiting probabilities in the three-candidate case for some minor cases left over from (Kamwa et al., 2023) under the impartial anonymous culture model.
Conflict of Interest and Funding Statement: The authors have no relevant financial or non-financial interests to disclose. The authors did not receive support from any organization for the submitted work.
Data Availability Statement: Most of the data used in this article is publicly available at (Mattei & Walsh, 2013), (Otis, 2022), and https://github.com/mggg/scot-elex. The rest is available upon request.
Appendix
We provide proofs of Proposition 3 and Proposition 4. The proofs use geometric techniques developed in (Saari & Tataru, 1999) and (Kamwa & Merlin, 2015).
Proof of Proposition 3 (i)
@@ 6 Conclusion
Step 1 First, we find the probability that $A$ and $B$ are the plurality and IRV winners respectively. Following the arguments of (Saari & Tataru, 1999), the probability that (1) and (3) are satisfied is equal to the area of the spherical simplex $S$ defined by these inequalities on the surface of the unit sphere in $\mathbb{R}^{3}$, divided by the area of this sphere. Let
@@
$\alpha_{13}$ $=\angle(\mathbf{v}_{1},\mathbf{v}_{3})=\cos^{-1}(\frac{-\mathbf{v}_{1}\cdot\mathbf{v}_{3}}{\lVert\mathbf{v}_{1}\rVert\lVert\mathbf{v}_{3}\rVert})=\cos^{-1}(\frac{4}{2\cdot\sqrt{6}})=\cos^{-1}(\frac{2}{\sqrt{6}})$
@@
Step 2: Next, we find the probability that additionally (5) is satisfied. Following (Kamwa & Merlin, 2015), we introduce the modified inequality with parameter $t\in[0,1],$
$a_{1}+a_{2}>b_{1}+b_{2}+t(b_{2}-b_{1}).$ (9)
@@
$dVol_{n}=\frac{1}{n-1}\sum_{1\leq j<k\leq n}Vol_{n-2}(S_{j}\cap S_{k})d\alpha_{jk}$
@@
$Vol(S_{1}\cap S_{4})$ $=\angle(P_{124},P_{134})=\cos^{-1}(\frac{0}{\sqrt{2}\cdot 2\sqrt{3}})=\frac{\pi}{2}$
$Vol(S_{2}\cap S_{4})$ $=\angle(P_{124},P_{234})=\cos^{-1}(\frac{\sqrt{3}t}{\sqrt{5t^{2}+3}})$
$Vol(S_{3}\cap S_{4})$ $=\angle(P_{134},P_{234})=\cos^{-1}(\frac{t}{\sqrt{2}\sqrt{5t^{2}+3}})$
@@
$\frac{1}{2\pi^{2}}\left[0.04077671(2\pi^{2})+\frac{1}{2}(-0.99679+0.11231+0.312545)\right]=0.026289.$
@@
Since the choice of $A$ and $B$ as plurality and IRV winners was arbitrary, the probability of a negative participation paradox is $6\times 0.026289\approx 0.157737$
@@
Proof of (ii) Since a negative participation paradox can only occur when the IRV and plurality winners are distinct, the probability that an election demonstrates a negative participation paradox given that these winners are different is equal to the probability of negative participation paradox divided by the probability that the IRV and plurality winners are distinct. The latter was determined in Step 1; hence this probability is equal to $0.026289/0.04077671\approx 0.644717$
@@
Step 2: Next, we find the probability that additionally (7) is satisfied. We introduce the inequality $a_{0}+a_{1}+a_{2}>b_{0}+b_{1}+b_{2}+t(b_{2}-b_{1})$. When $t=0$ this reduces to a (2); when $t=1$ we get the required extra condition.
@@
$Vol(S_{1}\cap S_{4})$ $=\angle(P_{124},P_{134})=\cos^{-1}(\frac{0}{\sqrt{2}\cdot 2\sqrt{18}})=\frac{\pi}{2}$
$Vol(S_{2}\cap S_{4})$ $=\angle(P_{124},P_{234})=\cos^{-1}(\frac{18t}{\sqrt{2}\cdot 3\sqrt{2(13t^{2}+9)}})=\cos^{-1}(\frac{3t}{\sqrt{13t^{2}+9}})$
$Vol(S_{3}\cap S_{4})$ $=\angle(P_{134},P_{234})=\cos^{-1}(\frac{18t}{3\sqrt{2}\cdot 3\sqrt{2(13t^{2}+9)}})=\cos^{-1}(\frac{t}{\sqrt{13t^{2}+9}})$
@@
Proof of (ii) As with the case with complete ballots, the probability that an election demonstrates a negative participation paradox given that these winners are different is equal to the probability of negative participation paradox divided by the probability that the IRV and plurality winners are distinct. The latter was determined in Step 1; hence this probability is equal to $0.023976912/0.03563737\approx 0.672803.\qed$